OpenStudy (anonymous):

3x^2+x=4 solve equation by method of choice (QUAD FORMULA)

hero (hero):

3x^2 + x - 4 = 0 I'd start off with that

OpenStudy (mimi_x3):

3x^2 + x - 4 = 0 (3x-3)(3x+4) /3 = 0 (x-1)(3x+4) x = 1 , -3/4

OpenStudy (mimi_x3):

my bad x = 1 , -4/3

hero (hero):

3x^2 + 4x - 3x - 4 = 0 x(3x+ 4)-1(3x+4) = 0 (3x+4)(x-1) = 0 x = 1, -4/3